# Keep It Real: Challenges in Attacking Compression-Based Adversarial Purification

### v1

@@ Abstract
The first principle is that you must not fool yourself–and you are the easiest person to fool.
@@ 1 Introduction
Lossy image compression often discards details deemed perceptually unimportant, which may include the subtle perturbations introduced by adversarial attacks. Early work argued that standard codecs like JPEG yield modest robustness gains [Guo et al., 2017], though they often introduce compression artifacts that push images outside the classifier’s training distribution. More recently, learned compression models have promised to bridge this gap by generating visually plausible reconstructions that may also remove stronger adversarial noise [Räber et al., 2025].
However, many proposed defenses—especially preprocessing-based ones—have been criticized for relying on gradient masking; they make general gradient-based attacks harder without genuinely increasing robustness [Shin and Song, 2017]. When attackers adapt to circumvent gradient obfuscation (e.g., by approximating or smoothing gradients), many defenses fail [Carlini et al., 2019, Tramer et al., 2020, Sheatsley et al., 2023, Zhang et al., 2025]. This raises two critical questions:
@@
In this paper, we systematically evaluate the robustness of compression-based defenses and isolate the role of realism. Our analysis builds on the findings in [Räber et al., 2025], where it was argued that human-aligned compression contributes to robustness. We re-evaluate their claims under rigorous adversarial threat models and identify shortcomings in their evaluation protocol. We show that their observed robustness stemmed not merely from compression, but specifically from realism. We reinforce this conclusion through extensive evaluation across a broader and newer set of learned compression models.
@@ 2.1 Adversarial Robustness
Shortly after the success of AlexNet [Krizhevsky et al., 2012], it was found that neural networks are very susceptible to adversarial attacks [Szegedy et al., 2014, Goodfellow et al., 2015]. Here, an adversary adds (usually) small and imperceptible perturbations to an image such that a model mislabels it.
@@
Many attacks have been developed over the years with various benefits and drawbacks. Some of the most noteworthy are FGSM (Fast Gradient Sign Method) [Goodfellow et al., 2015], iterated FGSM or iFGSM [Kurakin et al., 2017], CW [Carlini and Wagner, 2017a], and PGD (Projected Gradient Descent) [Madry et al., 2019]. These attacks fall into two categories: FGSM, iFGSM, and PGD are $l_{\infty}$-bounded attacks, and CW is an $l_{2}$-bounded attack. We refer the reader to the original papers for specific details, but include some details in Appendix˜C. The TLDR is: PGD has a perturbation budget $\epsilon$ and an iteration budget $n$. PGD does projected gradient descent for $n$ iterations to find an adversarial example at most $\epsilon$ distance in $l_{\infty}$ from the original image.
When the defense includes randomness, a common attack augmentation is EoT (Expectation over Transformation), where gradients are averaged over multiple backward passes [Athalye et al., 2018a, Zhang et al., 2025]. If the defense causes gradient masking, Athalye et al. [2018a] suggested Backward Pass Differentiable Approximation (BPDA), where the defense is used during the forward pass, but a differentiable approximation is used during the backward pass. The simplest is to use the identity function, while more advanced methods might train a differentiable surrogate model $g^{\prime}$ that emulates the defense $g$; i.e., $g(x)\approx g^{\prime}(x)$.
Lastly, often the best attack results are found using adaptive attacks [Carlini et al., 2019, Tramer et al., 2020, Sheatsley et al., 2023, Zhang et al., 2025]. Here, the attacks, for instance, the optimization objective, are adjusted to the defense. This work focuses on adaptive attacks with PGD as the underlying optimization method.
@@
Given the prevalence of adversarial attacks, many researchers have explored how to defend against them. We broadly view this in three groups: Architecture improvements [Singh et al., 2023, Wu et al., 2023, Fort and Lakshminarayanan, 2024], adversarial training [Szegedy et al., 2014, Goodfellow et al., 2015, Madry et al., 2019, Pang et al., 2020a, Fort and Lakshminarayanan, 2024], and adversarial purification [Nie et al., 2022, Frosio and Kautz, 2023, Zollicoffer et al., 2025, Räber et al., 2025]. The architecture branch focuses on making the models more robust by design, for instance, by taking inspiration from biology to mimic the human eye when training a ResNet model [Fort and Lakshminarayanan, 2024]. Adversarial training consists of showing adversarial examples during training to ensure correct classifications. This has yielded positive results [Szegedy et al., 2014, Madry et al., 2019], but models can be broken [Zhang et al., 2025].
@@
Diffusion models have been proposed as a way to remove adversarial noise from input images [Nie et al., 2022, Lee and Kim, 2023]. These approaches are conceptually similar to compression-based defenses with realism: both aim to project adversarial examples back onto the manifold of natural images to restore classifier performance. However, prior work on diffusion-based purification has not explicitly investigated the role of realism as a contributing factor to robustness. One major drawback of diffusion-based defenses is their computational cost [Dhariwal and Nichol, 2021, Yang and Mandt, 2023a]. But simply making gradients difficult to compute does not equate to genuine robustness. The purification method proposed by Nie et al. [2022] was later defeated by Lee and Kim [2023]. However, the latter only evaluated their improved defense under the attack that broke the former. They did not develop new adaptive attacks tailored to their defense—a strategy that past research suggests would likely reduce the effectiveness of the defense. The history of adversarial robustness research shows that defenses that are not tested under strong, tailored attacks often overstate their robustness [Athalye et al., 2018a, Carlini et al., 2019, Tramer et al., 2020].
@@ 2.2 Realism in Image Compression
However, metrics such as PSNR, MS-SSIM [Wang et al., 2003], or LPIPS [Zhang et al., 2018] correlate poorly with human perception, and directly quantifying perceptual distortion remains a challenging problem. Instead of a perceptual distortion metric, one can measure both distortion and realism and optimize the compression model for both. Realism can be formally defined as:
@@
where $d(\cdot,\cdot)$ is a divergence measure, such as Kullback-Leibler. Unlike distortion, realism is a no-reference metric, requiring only the generated image distribution to match that of natural images. Although measuring realism remains challenging [Theis, 2024], a widely used proxy is the Fréchet Inception Distance (FID) [Heusel et al., 2017], which compares the distributions extracted by a pretrained Inception network. FID has gained popularity for capturing both fidelity and diversity in generated samples.
@@
Compression models can be broadly categorized into those optimized for distortion (e.g., JPEG [Wallace, 1991], Hyperprior [Ballé et al., 2016], ELIC [He et al., 2022a]) and those explicitly trained to maximize realism (e.g., HiFiC [Mentzer et al., 2020], PO-ELIC [He et al., 2022b]). More recent approaches, such as MRIC [Agustsson et al., 2023] and CRDR [Iwai et al., 2024], provide explicit control over the realism–distortion tradeoff within a single network by conditioning on $\lambda$ and $\beta$. This controllability enables direct investigation of how realism influences adversarial robustness.
@@ 2.3 Compression as Adversarial Defense
Using compression as a defense for neural networks is not a new idea. It has been explored for almost a decade [Guo et al., 2017, Dziugaite et al., 2016, Das et al., 2017, Jia et al., 2019, Ferrari et al., 2023] where some authors also explored using iterated compression and decompression cycles [Ferrari et al., 2023, Räber et al., 2025]. Two key works in this area are by Guo et al. [2017] and Shin and Song [2017]; the former argued that JPEG compression as a preprocessing step is a very effective adversarial defense while the latter showed that, by making JPEG differentiable, this defense could be bypassed entirely–highlighting the necessity of properly evaluating a defense.
@@ 3 Methodology
Our work evaluates the compression models’ robustness; thus, we focus on the ImageNet classification benchmark, a well-established benchmark with high-resolution images. This gives us many options for pretrained models, allowing a more exhaustive model evaluation. If otherwise not stated, we use the full validation split of the ImageNet dataset (50000 images). We used two different classification models, ResNet50 [He et al., 2016] and ViT B 16 [Dosovitskiy et al., 2020]. For PyTorch models, we use the improved weights IMAGENET1K_V2 for ResNet50 and the default weights IMAGENET1K_V1 for ViT B 16. For TensorFlow models, we use the default ResNet50 pretrained weights imagenet. Certain ablations are only done for ResNet to reduce the compute load; the results have only minor differences to ViT.
@@ 3.1 Defenses
As we demonstrate later, realism plays a crucial role in the robustness of this pipeline. Therefore, we focus our experiments on models that either explicitly control the level of realism or exist in both standard and enhanced realism variants. In particular, MRIC [Agustsson et al., 2023] and CRDR [Iwai et al., 2024] offer variable realism settings, and we denote their low- and high-realism variants as MRIC LR, MRIC HR, CRDR LR, and CRDR HR, respectively. We also consider models available in both standard rate-distortion and rate-distortion-realism versions, such as Hyperprior [Ballé et al., 2018] and HiFiC [Mentzer et al., 2020]. To further validate the importance of realism, we include JPEG [Wallace, 1991] and ELIC [He et al., 2022a] (noting that the high-realism version, PO-ELIC [He et al., 2022b], does not have publicly available pretrained weights), and show that these consistently underperform compared to high-realism compression models. For white-box attacks, we require access to gradients; for learned compression models, gradient computation is natively supported, while for JPEG, we employ a continuous relaxation approach [Shin and Song, 2017]. We focus on VAE-based methods as diffusion or INR-based approaches are prohibitively expensive to evaluate at scale [Nie et al., 2022, Lee and Kim, 2023].
@@ 3.2 Threat models
We define our own threat models that encompass the prior experiments from Räber et al. [2025] to ensure proper evaluation [Carlini and Wagner, 2017b, Carlini et al., 2019, Rando et al., 2025]. The threat models follow standard formulations from Biggio et al. [2013], Carlini and Wagner [2017b], Pang et al. [2020b]. Specifically, we consider $l_{\infty}$ untargeted attacks with perturbation budget $\epsilon$, i.e., for an original image $x$ and perturbed image $x^{\prime}$, we have $\|x-x^{\prime}\|_{\infty}\leq\epsilon$.
Our primary focus is on PGD attacks, widely regarded as among the strongest given sufficient objective formulation and computational resources (see Appendix˜C). We also include adaptive attacks, where the adversary, aware of the defense mechanism, tailors the attack accordingly. Adaptation in this context means the adversary can specialize the attack by studying the defense and looking for ways to defeat it [Tramer et al., 2020]. These adaptive attacks still use PGD (see Section˜3.3 for details). If otherwise not stated, we use 10 PGD iterations.
Lastly, we consider three adversary knowledge levels as described in Carlini and Wagner [2017b]. Black-box (BB): The adversary does not know the defense or its existence. The attacker can create adversarial perturbations with respect to the gradients of the classifier or the outputs of the classifier. Gray-box (GB): The adversary knows the defense is present and can use it for the forward pass only; they cannot compute gradients through the defense. This threat variant is used for adaptive attacks. White-box (WB): The adversary has complete knowledge and access to the defense. The gradients of the defense and the output of the combined defense and classifier pipeline are used in the attack.
@@ 3.3 Adaptive Attacks Against Compression Defenses
A common pitfall in adversarial robustness papers is the lack of honest effort in implementing proper attacks against one’s own defense [Carlini et al., 2019, Tramer et al., 2020, Sheatsley et al., 2023]. In our evaluation, we conduct adaptive attacks on all compression models, which means the attacks are tailored to exploit the target model’s weaknesses. This section reviews the adaptive attacks used against the compression defenses. We use the following notation: $f$ is the image classifier, $g$ is a compression defense (compression and decompression), and $h=f\circ g$. $x$ is an image with label $y$, and $L$ is a loss function. We use cross-entropy as the loss function for all our experiments (except ACM). $\nabla_{x}f$ is the gradient of $L(f(x),y)$ with respect to $x$.
@@
In the second adaptive attack, we use BPDA with a standard U-Net [Ronneberger et al., 2015] trained to approximate the compression defense $g$ (see training details in Appendix˜F). The idea is that if $g$ primarily causes gradient masking, then replacing it with a differentiable proxy $g^{\prime}$ should yield meaningful gradients for the attack. During the forward pass, we use the actual defense, $h(x)=f(g(x))$, while the backward pass substitutes $g$ with $g^{\prime}$, yielding gradients $\nabla_{x}h\coloneq\nabla_{x}(f\circ g^{\prime})(x)$.
@@
All the VAE-based compression defenses are deterministic; thus, EoT [Athalye et al., 2018b] should not be required for these defenses [Athalye et al., 2018a].
@@ 4.1 Role of Realism
To isolate the impact of realism, we analyze two compression models where realism can be explicitly controlled. As shown in Figure˜6 (Section˜G.2), increasing realism in reconstructed images monotonically improves robustness. While distortion has been the focus of extensive prior work [Dziugaite et al., 2016, Shin and Song, 2017, Räber et al., 2025], realism remains largely unexplored as a factor in defense performance. Our results highlight a key difference: distortion presents an inherent trade-off. If the distortion is too low, adversarial noise is preserved. If the distortion is too high, the reconstruction discards critical semantic information, making the image unrecognizable. In both cases, the defense fails—either by retaining harmful perturbations or by producing images that are implausible under the natural data distribution. In contrast, increasing realism consistently improves robustness without the trade-offs typically associated with distortion. Prior to this work, the role of realism in adversarial defenses had not been thoroughly investigated.
@@ 4.4 Comparison to Diffusion-Based Defenses
Table 4: Robust accuracy under PGD 4/255 attack running for ten iterations for diffusion-based defenses and the CRDR compression-based defense. The robust accuracy numbers of the diffusion-based models are taken from Lee and Kim [2023].
@@ 4.5 Highlight of Extra Results in the Appendix
Instead of evaluating against standard (adversarially weak) classifiers, one can use pretrained robust classifiers to assess whether realism offers additional benefits in already robust settings. We ran the compression-based defense on the top 11 models from RobustBench [Croce et al., 2021, Singh et al., 2023, Bai et al., 2024, Amini et al., 2024, Xu et al., 2025, Liu et al., 2025]. Our RobustBench results can be seen in Table˜11 in Section˜G.3. The results show that applying CRDR with high quality and realism does not improve performance—in fact, it consistently reduces accuracy across all tested robust models.
Iterative compression as a defense has shown promising results [Räber et al., 2025], but its robustness largely stems from gradient masking. When attacking the defense using fewer iterations, robust accuracy drops significantly, revealing the underlying vulnerability once gradient masking is circumvented. See Section˜G.4 for detailed results.

### v3

# Reliable ML from Unreliable Data

@@ Abstract
The first principle is that you must not fool yourself–
   and you are the easiest person to fool.
@@ 1 Introduction
Lossy image compression often discards details deemed perceptually unimportant, which may include the subtle perturbations introduced by adversarial attacks. Early work argued that standard codecs like JPEG yield modest robustness gains [1], though they often introduce compression artifacts that push images outside the classifier’s training distribution. More recently, learned compression models have promised to bridge this gap by generating visually plausible reconstructions that may also remove stronger adversarial noise [2].
However, many proposed defenses—especially preprocessing-based ones—have been criticized for relying on gradient masking; they make general gradient-based attacks harder without genuinely increasing robustness [3]. When attackers adapt to circumvent gradient obfuscation (e.g., by approximating or smoothing gradients), many defenses fail [4, 5, 6, 7]. This raises two critical questions:
@@
In this paper, we systematically evaluate the robustness of compression-based defenses and isolate the role of realism. Our analysis builds on the findings in [2], where it was argued that human-aligned compression contributes to robustness. We re-evaluate their claims under rigorous adversarial threat models and identify shortcomings in their evaluation protocol. We show that their observed robustness stemmed not merely from compression, but specifically from realism. We reinforce this conclusion through extensive evaluation across a broader and newer set of learned compression models.
@@ 2.1 Adversarial Robustness
Shortly after the success of AlexNet [8], it was found that neural networks are very susceptible to adversarial attacks [9, 10]. Here, an adversary adds (usually) small and imperceptible perturbations to an image such that a model mislabels it.
@@
Many attacks have been developed over the years with various benefits and drawbacks. Some of the most noteworthy are FGSM (Fast Gradient Sign Method) [10], iterated FGSM or iFGSM [11], CW [12], and PGD (Projected Gradient Descent) [13]. These attacks fall into two categories: FGSM, iFGSM, and PGD are $l_{\infty}$-bounded attacks, and CW is an $l_{2}$-bounded attack. We refer the reader to the original papers for specific details, but include some details in Appendix˜C. The TLDR is: PGD has a perturbation budget $\epsilon$ and an iteration budget $n$. PGD does projected gradient descent for $n$ iterations to find an adversarial example at most $\epsilon$ distance in $l_{\infty}$ from the original image.
When the defense includes randomness, a common attack augmentation is EoT (Expectation over Transformation), where gradients are averaged over multiple backward passes [14, 7]. If the defense causes gradient masking, Athalye et al. [14] suggested Backward Pass Differentiable Approximation (BPDA), where the defense is used during the forward pass, but a differentiable approximation is used during the backward pass. The simplest is to use the identity function, while more advanced methods might train a differentiable surrogate model $g^{\prime}$ that emulates the defense $g$; i.e., $g(x)\approx g^{\prime}(x)$.
Lastly, often the best attack results are found using adaptive attacks [4, 5, 6, 7]. Here, the attacks, for instance, the optimization objective, are adjusted to the defense. This work focuses on adaptive attacks with PGD as the underlying optimization method.
@@
Given the prevalence of adversarial attacks, many researchers have explored how to defend against them. We broadly view this in three groups: Architecture improvements [15, 16, 17], adversarial training [9, 10, 13, 18, 17], and adversarial purification [19, 20, 21, 2]. The architecture branch focuses on making the models more robust by design, for instance, by taking inspiration from biology to mimic the human eye when training a ResNet model [17]. Adversarial training consists of showing adversarial examples during training to ensure correct classifications. This has yielded positive results [9, 13], but models can be broken [7].
@@
Diffusion models have been proposed as a way to remove adversarial noise from input images [19, 23]. These approaches are conceptually similar to compression-based defenses with realism: both aim to project adversarial examples back onto the manifold of natural images to restore classifier performance. However, prior work on diffusion-based purification has not explicitly investigated the role of realism as a contributing factor to robustness. One major drawback of diffusion-based defenses is their computational cost [24, 25]. But simply making gradients difficult to compute does not equate to genuine robustness. The purification method proposed by Nie et al. [19] was later defeated by Lee and Kim [23]. However, the latter only evaluated their improved defense under the attack that broke the former. They did not develop new adaptive attacks tailored to their defense—a strategy that past research suggests would likely reduce the effectiveness of the defense. The history of adversarial robustness research shows that defenses that are not tested under strong, tailored attacks often overstate their robustness [14, 4, 5].
@@ 2.2 Realism in Image Compression
However, metrics such as PSNR, MS-SSIM [26], or LPIPS [27] correlate poorly with human perception, and directly quantifying perceptual distortion remains a challenging problem. Instead of a perceptual distortion metric, one can measure both distortion and realism and optimize the compression model for both. Realism can be formally defined as:
@@
where $d(\cdot,\cdot)$ is a divergence measure, such as Kullback-Leibler. Unlike distortion, realism is a no-reference metric, requiring only the generated image distribution to match that of natural images. Although measuring realism remains challenging [28], a widely used proxy is the Fréchet Inception Distance (FID) [29], which compares the distributions extracted by a pretrained Inception network. FID has gained popularity for capturing both fidelity and diversity in generated samples.
@@
Compression models can be broadly categorized into those optimized for distortion (e.g., JPEG [30], Hyperprior [31], ELIC [32]) and those explicitly trained to maximize realism (e.g., HiFiC [33], PO-ELIC [34], WD [35], ConHa [36]). More recent approaches, such as MRIC [37] and CRDR [38], provide control over the realism–distortion tradeoff within a single network by conditioning on $\lambda$ and $\beta$. This controllability enables direct investigation of how realism influences adversarial robustness.
@@ 2.3 Compression as Adversarial Defense
Using compression as a defense for neural networks is not a new idea. It has been explored for almost a decade [1, 39, 40, 41, 42] where some authors also explored using iterated compression and decompression cycles [42, 2]. Two key works in this area are by Guo et al. [1] and Shin and Song [3]; the former argued that JPEG compression as a preprocessing step is a very effective adversarial defense while the latter showed that, by making JPEG differentiable, this defense could be bypassed entirely–highlighting the necessity of properly evaluating a defense.
@@ 3 Methodology
Our work evaluates the compression models’ robustness; thus, we focus on the ImageNet classification benchmark, a well-established benchmark with high-resolution images. This gives us many options for pretrained models, allowing a more exhaustive model evaluation. If otherwise not stated, we use the full validation split of the ImageNet dataset (50000 images). We used two different classification models, ResNet50 [43] and ViT B 16 [44]. For PyTorch models, we use the improved weights IMAGENET1K_V2 for ResNet50 and the default weights IMAGENET1K_V1 for ViT B 16. For TensorFlow models, we use the default ResNet50 pretrained weights imagenet. Certain ablations are only done for ResNet to reduce the compute load; the results have only minor differences to ViT.
@@ 3.1 Defenses
As we demonstrate later, realism plays a crucial role in the robustness of this pipeline. Therefore, we focus our experiments on models that either explicitly control the level of realism or exist in both standard and enhanced realism variants. In particular, MRIC [37] and CRDR [38] offer variable realism settings, and we denote their low- and high-realism variants as MRIC LR, MRIC HR, CRDR LR, and CRDR HR, respectively. We also consider models available in both standard rate-distortion and rate-distortion-realism versions, such as Hyperprior [45] and HiFiC [33]. To further validate the importance of realism, we include JPEG [30] and ELIC [32] (noting that the high-realism version, PO-ELIC [34], does not have publicly available pretrained weights), and show that these consistently underperform compared to high-realism compression models. For white-box attacks, we require access to gradients; for learned compression models, gradient computation is natively supported, while for JPEG, we employ a continuous relaxation approach [3]. We focus on VAE-based methods as diffusion or INR-based approaches are prohibitively expensive to evaluate at scale [19, 23].
@@ 3.2 Threat models
We define our own threat models that encompass the prior experiments from Räber et al. [2] to ensure proper evaluation [46, 4, 47]. The threat models follow standard formulations from Biggio et al. [48], Carlini and Wagner [46], Pang et al. [49]. Specifically, we consider $l_{\infty}$ untargeted attacks with perturbation budget $\epsilon$, i.e., for an original image $x$ and perturbed image $x^{\prime}$, we have $\|x-x^{\prime}\|_{\infty}\leq\epsilon$.
Our primary focus is on PGD attacks, widely regarded as among the strongest given sufficient objective formulation and computational resources (see Appendix˜C). We also include adaptive attacks, where the adversary, aware of the defense mechanism, tailors the attack accordingly. Adaptation in this context means the adversary can specialize the attack by studying the defense and looking for ways to defeat it [5]. These adaptive attacks still use PGD (see Section˜3.3 for details). If otherwise not stated, we use 10 PGD iterations.
Lastly, we consider three adversary knowledge levels as described in Carlini and Wagner [46]. Black-box (BB): The adversary does not know the defense or its existence. The attacker can create adversarial perturbations with respect to the gradients of the classifier or the outputs of the classifier. Gray-box (GB): The adversary knows the defense is present and can use it for the forward pass only; they cannot compute gradients through the defense. This threat variant is used for adaptive attacks. White-box (WB): The adversary has complete knowledge and access to the defense. The gradients of the defense and the output of the combined defense and classifier pipeline are used in the attack.
@@ 3.3 Adaptive Attacks Against Compression Defenses
A common pitfall in adversarial robustness papers is the lack of honest effort in implementing proper attacks against one’s own defense [4, 5, 6]. In our evaluation, we conduct adaptive attacks on all compression models, which means the attacks are tailored to exploit the target model’s weaknesses. This section reviews the adaptive attacks used against the compression defenses. We use the following notation: $f$ is the image classifier, $g$ is a compression defense (compression and decompression), and $h=f\circ g$. $x$ is an image with label $y$, and $L$ is a loss function. We use cross-entropy as the loss function for all our experiments (except ACM). $\nabla_{x}f$ is the gradient of $f(x)$ with respect to $x$.
@@
In the second adaptive attack, we use BPDA with a standard U-Net [50] trained to approximate the compression defense $g$ (see training details in Appendix˜F). The idea is that if $g$ primarily causes gradient masking, then replacing it with a differentiable proxy $g^{\prime}$ should yield meaningful gradients for the attack. During the forward pass, we use the actual defense, $h(x)=f(g(x))$, while the backward pass substitutes $g$ with $g^{\prime}$, yielding gradients $\nabla_{x}h\coloneq\nabla_{x}(f\circ g^{\prime})(x)$.
@@
All the VAE-based compression defenses are deterministic; thus, EoT [51] should not be required for these defenses [14].
@@ 4.1 Role of Realism
To isolate the impact of realism, we analyze two compression models where realism can be explicitly controlled. As shown in Figure˜6 (Section˜G.2), increasing realism in reconstructed images monotonically improves robustness. While distortion has been the focus of extensive prior work [39, 3, 2], realism remains largely unexplored as a factor in defense performance. Our results highlight a key difference: distortion presents an inherent trade-off. If the distortion is too low, adversarial noise is preserved. If the distortion is too high, the reconstruction discards critical semantic information, making the image unrecognizable. In both cases, the defense fails—either by retaining harmful perturbations or by producing images that are implausible under the natural data distribution. In contrast, increasing realism consistently improves robustness without the trade-offs typically associated with distortion. Prior to this work, the role of realism in adversarial defenses had not been thoroughly investigated.
@@ 4.4 Comparison to Diffusion-Based Defenses
Table 4: Robust accuracy under PGD 4/255 attack running for ten iterations for diffusion-based defenses and the CRDR compression-based defense. The robust accuracy numbers of the diffusion-based models are taken from [23].
@@ 4.5 Highlight of Extra Results in the Appendix
Instead of evaluating against standard (adversarially weak) classifiers, one can use pretrained robust classifiers to assess whether realism offers additional benefits in already robust settings. We ran the compression-based defense on the top 11 models from RobustBench [54, 15, 55, 56, 57, 58]. Our RobustBench results can be seen in Table˜11 in Section˜G.3. The results show that applying CRDR with high quality and realism does not improve performance—in fact, it consistently reduces accuracy across all tested robust models.
Iterative compression as a defense has shown promising results [2], but its robustness largely stems from gradient masking. When attacking the defense using fewer iterations, robust accuracy drops significantly, revealing the underlying vulnerability once gradient masking is circumvented. See Section˜G.4 for detailed results.